OpenStudy (anonymous):

Anyone love doing math and wants to help me out??? WILL GIVE MEDAL AND FAN!!!

OpenStudy (anonymous):

um depends what it iss

OpenStudy (anonymous):

Algebra math :/ its summer school and there is a reason im there. Im not good at it :(

OpenStudy (mathmale):

I'd love to help you if I'm familiar with the type of problem you need help with. It's important that you please post your question up front, so that interested and qualified helpers may respond meaningfully. Please ditch the "medal and fan" business; it's not necessary.

OpenStudy (anonymous):

I'll help

OpenStudy (anonymous):

@mathmale you got it(: @vlltimelow I'll post the question.

OpenStudy (anonymous):

I could try

OpenStudy (anonymous):

Solve the given equation 1. -17 = 10x + 3 2. 9x + 5 = 7x + 3x - 2 3. 5(x-3) + 4x - 1= 2x + 3(x-2) Determine whether the equation is an identity, a conditional equation or an inconsistent equation. 4. 8x - 3 = 4x + 4(x-1) 5. 5/4x + 1/2 = 2x - 1/2 6. 9(x-3) = 9x - 27

OpenStudy (mathmale):

You've gone from no questions to six questions just like that! One at a time, please. -17 = 10x + 3 is to be solved for x. To accomplish this, subtract 3 from both sides of the equation (to isolate 10x).

OpenStudy (anonymous):

Okay, let me work it out

OpenStudy (anonymous):

Okay, I got x = 2??

OpenStudy (precal):

no its not 2

OpenStudy (precal):

-17 = 10x + 3 -3 -3 -20=10x divide both sides by 10 -2=x

OpenStudy (anonymous):

Ohhhh -2?

OpenStudy (anonymous):

Okay, im gonna work out the next one

OpenStudy (anonymous):

Umm is it x = 7 or x = -3 ?? idk

OpenStudy (anonymous):

x=7

OpenStudy (precal):

9x + 5 = 7x + 3x - 2 9x+5=10x-2 -9x -9x 5=1x-2 +2 +2 7=x

OpenStudy (anonymous):

How do you solve number 3?

OpenStudy (anonymous):

So I'm a bit late but any multiplication with different symbols (+ plus -) will equal a negative (-) and any multiplication with the same symbols will equal a positive (+)

OpenStudy (anonymous):

@FreeWilly thanks(:

OpenStudy (precal):

you need to distribute whatever is in front of the parenthesis example 2(x+4) is 2x+8

OpenStudy (anonymous):



OpenStudy (precal):



OpenStudy (precal):

I hope you realize I just made that example up

OpenStudy (anonymous):

Cool?